Two sinusoidal waves travelling in the same direction with the same amplitude, wavelength, and speed, interfere with each other to give the resultant wave: y_res (x,t) = 2 cm sin(4πx-60πt+π/3). The amplitude of the individual waves generating this wave is:


8 cm


2/√3 cm


4/√3 cm


4 cm


2 cm


1
Expert's answer
2021-05-23T18:54:09-0400

Gives

"Y=2sin(4\pi x-60\pi t+\frac{\pi}{3})"

Wave is same amplitude speed wavelength

"A'^2=A_1^2+A_2^2+2A_1A_2cos\phi"

"\phi=\frac{\pi}{3}"

Let us Assume same amplitude is "A_1=A_2=" A

"A'^2=A^2+A^2+2A^2cos(\frac{\pi}{3})"

"2^2=A^2+A^2+2A^2cos(\frac{\pi}{3})"

"4=2A^2+A^2=3A^2"

"3A^2=4"

"A=\sqrt{\frac{4}{3}}"

"A=\frac{2}{\sqrt3}"